A student carried out an experiment trying to prepare tetra-methyl-ammonium tri-iodide.

The student weighed 0.94 g tetra-methyl-ammonium iodide into a beaker containing 25 mL ethanol, and then added 1.22 g of iodine. The mixture was heated gently, whilst stirring, on a hotplate until tetra-methyl-ammonium iodide, a white powder, was completely dissolved. After the mixture was allowed to cool down to room temperature, the mixture was further cooled on ice. After filtration on a Buechner funnel, washing with diethyl ether and leaving it under vacuum suction for a couple of minutes, the student recovered the product and weighed it. The mass of the obtained tetra-methyl-ammonium triiodide was 0.21 g.

What is the yield of the reaction in percent [%]?



show your work please


1
Expert's answer
2021-11-18T22:02:02-0500

Molar mass of "C_4H_{12}IN" is 201.14


Molar mass of "C_4H_{12}I_3N" is 484.56



Moles of "C_4H_{12}IN=\frac{0.94}{201.14}=0.0047moles"


Molarity "= \frac{0.0047}{0.025}=0.188M"


Moles of "C_4H_{12}I_3N=\frac{0.21}{484.56}=0.00043moles"